### 25 Mar 2026 v1

# Mpemba effect in a two-dimensional bistable potential

Abstract
We present an exactly solvable model of the Mpemba effect in an overdamped Langevin system confined in a two-dimensional radially symmetric bistable potential. The potential is constructed as a piecewise quadratic-logarithmic function that is continuous and differentiable at the matching radii, enabling an exact mapping of the corresponding Fokker-Planck operator to a Schrödinger-type eigenvalue problem. The relaxation spectrum and eigenmodes are obtained analytically in each region in terms of confluent hypergeometric functions, with eigenvalues determined from matching conditions. Focusing on isotropic equilibrium initial states at inverse temperature $\beta_{\rm ini}$ quenched to a bath at inverse temperature $\beta$, we derive explicit expressions for the mode amplitudes governing long-time relaxation. We demonstrate that the coefficient of the slowest mode exhibits non-monotonic dependence on $\beta_{\rm ini}$ and identify a sufficient crossing condition for the Kullback-Leibler divergence in terms of the two slowest modes, if the global minimum of the potential is located far away from the origin and the second minimum exists near the origin. For corresponding parameters, we demonstrate that the Mpemba effect can be realized. Our results provide a rare example of an analytically tractable two-dimensional model exhibiting anomalous relaxation without any confining walls, extending previous one-dimensional constructions with a hard wall and clarifying the role of radial geometry in nonequilibrium relaxation phenomena.
I Introduction
The Mpemba effect refers to the anomalous relaxation phenomenon in which an initially hotter system, quenched into contact with a cold bath, reaches equilibrium faster than the same system prepared at a lower initial temperature. The effect was first reported by Mpemba and Osborne [1], who observed that hotter water can freeze faster than colder water. Although its very existence has been debated [2, 3], theoretical studies soon after the debates already suggested possible scenarios for faster freezing from higher temperatures [4, 5]. In recent years, numerous experimental and theoretical works have confirmed Mpemba-like processes in which a “temperature-like” observable exhibits a crossing, such that a system initialized at higher temperature subsequently relaxes faster than the same system started at lower temperature.
The Mpemba effect has been reported in a broad spectrum of classical systems, including colloidal particles in optical traps [6, 7], thermostated granular fluids [5, 8, 9, 10, 11, 12, 13], optical resonators [14, 15, 16], inertial suspensions [17, 18], spin glasses [19], molecular binary mixtures [20], phase transitions [21, 22], and a wide variety of Markovian model [4, 23, 24, 25, 26, 27, 28], non-Markovian models [29, 30] and active matters [31]. In addition, quantum analogues of the effect have been demonstrated in open quantum systems [32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49], establishing the Mpemba effect as a general nonequilibrium relaxation phenomenon.
Several reviews now provide unified perspectives on both classical and quantum Mpemba effects [50, 51, 52], emphasizing their connection to nonequilibrium speed limits and anomalous relaxation mechanisms. Two broad scenarios are typically distinguished [53, 54, 55, 56]: (i) comparing relaxation from an equilibrium versus a non-equilibrium initial condition, and (ii) comparing relaxation from two equilibrium states prepared at different initial temperatures. The latter scenario, which we follow in this paper, has led to general unifying frameworks [55, 56].
A central theoretical picture, proposed by Lu and Raz [4], is that the Mpemba effect arises in systems with a non-convex free-energy landscape. Starting from an initial distribution localized near a metastable minimum, the mean relaxation time across the barrier decreases with increasing initial temperature, leading to the observed crossing phenomenon. This interpretation naturally extends to one-particle dynamics in bistable potentials, as confirmed in experiments on optically trapped colloids [6, 7]. Variants such as the strong Mpemba effect [24] have also been studied, though their relation to the original freezing anomaly is more tenuous. If we are interested in relaxations from equilibrium initial conditions, the Mpemba effect arises when the projection onto the slowest mode is smaller for the initially hotter state, even if its total distance from equilibrium is larger.
So far, we have only results for the motion of a particle in a discrete-level model, piece-wise linear potentials, and square box potentials [4, 5, 6, 57, 27]. There are no analytical studies of the motion of a particle confined to a smooth bistable potential, even in one-dimensional (1D) settings. Experimentally and theoretically, one-dimensional systems differ fundamentally from higher-dimensional systems, such as two-dimensional (2D) systems. Quite recently, Liu et al. [53, 54] have analyzed the motion in 1D asymmetric bistable potentials and found that the metastable picture is not correct for the realization of the Mpemba effect, but a confining wall is necessary to observe the Mpemba effect. To understand the robustness of the Mpemba effect, it is natural to analyze a 2D system to clarify the differences and common features in 1D and 2D systems, including whether confining walls are necessary to observe the Mpemba effect even in 2D.
In this work, we develop an exactly solvable model of the Mpemba effect in a smooth 2D asymmetric bistable potential by using a connection of three potentials taking into account matching conditions at the connection points. While previous exact or semi-analytical approaches have mostly been restricted to 1D models, our extension to two dimensions provides a natural and richer generalization. The analysis makes use of the mapping from the Fokker-Planck equation to a Schrödinger-type eigenvalue problem, which enables us to determine the relaxation spectrum and eigenmodes exactly. We show that the Mpemba effect of the Kullback-Leibler (KL) divergence [58], which is the best monotone measure, i.e., decreases monotonically with time, (and its inverse) can be realized in this setting without the introduction of any confining walls, thus broadening the landscape of analytically tractable nonequilibrium phenomena. Unlike previous criteria based on specific observables, our condition is formulated directly in terms of the full probability distribution via the KL divergence.
The contents of this paper are as follows. In the next section, we explain the model and setup, including the introduction of a solvable potential, and the mapping onto the Schrödinger equation. In Sec. III shows the mode-analysis of the mapped Schrödinger equation to obtain the solution of Schrödinger equation. In Sec. IV, we demonstrate the occurrence of the Mpemba effect by using the two-mode approximation. In Sec. V, we discuss our results. In Sec. VI, we present the summary of our results. In Appendix A, we briefly summarize the properties of hypergeometric functions we use. In Appendix B, we present the condition of the crossing of the KL divergence. In Appendix C, we discuss the contribution of higher-order terms of mode analysis. In Appendix D, we present the detailed conditions to have a peak of the slowest eigenmode against the initial temperature, mathematically.
II Model and setup
II.1 Formulation
Figure 1: A schematic of a bistable circular symmetric potential.
Let us consider the motion of a particle at a position $\bm{r}$ in a two-dimensional system. Here, the particle moves under the influence of a bistable potential $V(r)$ with $r:=|\bm{r}|$ (see Fig. 1). We assume that the motion of a particle obeys the dimensionless Langevin equation as
$\dot{\bm{r}}=-\frac{\partial V(r)}{\partial\bm{r}}+\bm{\xi}(t),$ (1)
where $\dot{\bm{r}}:=d\bm{r}(t)/dt$, and the noise $\xi_{\alpha}$, which is $\alpha$ component of $\bm{\xi}$, satisfies
$\langle\xi_{\alpha}(t)\rangle=0,\quad\langle\xi_{\alpha}(t)\xi_{\beta}(t^{\prime})\rangle=2T\delta_{\alpha\beta}\delta(t-t^{\prime})$ (2)
with the bath temperature $T$.
The corresponding Fokker-Planck equation can be written as
$\frac{\partial P(\bm{r},t)}{\partial t}$ $=\mathbb{L}P(\bm{r},t),$ (3)
$\mathbb{L}:$ $=\bm{\nabla}\cdot\left(\frac{\bm{r}}{r}V^{\prime}(r)\right)+T\nabla^{2},$ (4)
where $V^{\prime}(r):=dV(r)/dr$. The analysis of the Fokker-Planck equation follows the textbook [59].
Equations (3) and (4) can be rewritten as
$\frac{\partial P(r,t)}{\partial t}$ $=-\bm{\nabla}\cdot\bm{J}(\bm{r},t),\quad\bm{J}(\bm{r},t):=\left\{T\bm{\nabla}+\frac{\bm{r}}{r}V^{\prime}(r)\right\}P(\bm{r},t).$ (5)
Note that $P(\bm{r},t)$ should be isotropic if we begin with an isotropic initial condition, i.e. $P(\bm{r},t)=P(r,t)$ because the dynamics is governed by a circular symmetric potential $V(r)$.
We consider a quench process from $T_{\mathrm{ini}}=\beta_{\mathrm{ini}}^{-1}$ to $T=\beta^{-1}$. We also assume that the initial and final states are at equilibrium as
$P_{\mathrm{ini}}(r)$ $=P_{\mathrm{eq}}(r,\beta_{\mathrm{ini}}):=\frac{e^{-\beta_{\mathrm{ini}}V(r)}}{Z(\beta_{\mathrm{ini}})},$ (6)
$P(r,t\to\infty)$ $=P_{\mathrm{eq}}(r,\beta),$ (7)
with $Z(\beta):=2\pi\int_{0}^{\infty}drre^{-\beta V(r)}$, and $P_{\mathrm{eq}}(r,\beta)$ satisfies
$\mathbb{L}P_{\mathrm{eq}}(r,\beta)=0.$ (8)
II.2 A choice of solvable potential
In general, it is hard to obtain the exact solution of the Fokker-Planck equation with a double-well potential. However, if we choose a specific form of the potential $V(r)$, we can solve the Fokker-Planck equation. Thus, we aim to construct an exactly solvable radially symmetric double-well potential $V(r)$ in two dimensions. The requirements are: 1) $V(r)$ is continuous and differentiable across its domains, 2) $V(r)$ has a minimum at $r=0$, a maximum at $r=\xi$, and a second minimum at $r=\alpha>\xi$.
We take $V(r)$ the piecewise quadratic-logarithmic form:
$V(r)=\begin{cases}V_{\rm in}(r)=\dfrac{k_{\rm in}}{2}r^{2}+C_{\rm in},&0\leq r<r_{-},\\[5.0pt] V_{\rm mid}(r)=\dfrac{k_{\rm mid}}{2}r^{2}+b_{\rm mid}\ln r+C_{\rm mid},&r_{-}\leq r<r_{+},\\[5.0pt] V_{\rm out}(r)=\dfrac{k_{\rm out}}{2}r^{2}+b_{\rm out}\ln r+C_{\rm out},&r\geq r_{+},\end{cases}$ (9)
where $r_{\pm}$ denote matching points, and $C_{\rm in},C_{\rm mid}$, and $C_{\rm out}$ are, respectively, constants to enforce continuity. As will be shown, the conditions $b_{\mathrm{mid}}\propto\xi^{2}$ and $b_{\mathrm{out}}\propto\alpha^{2}$ should be satisfied. Interestingly, this specific form leads to an exactly solvable Fokker-Planck equation.
The derivatives in each region lead to
$V^{\prime}_{\rm in}(r)$ $=k_{\rm in}r,\quad V^{\prime}_{\rm mid}(r)=k_{\rm mid}\left(r-\frac{\xi^{2}}{r}\right),\quad V^{\prime}_{\rm out}(r)=k_{\rm out}\left(r-\frac{\alpha^{2}}{r}\right).$ (10)
The second derivative of $V(r)$ leads to
$V^{\prime\prime}_{\rm in}(r=0)=k_{\rm in},\quad V^{\prime\prime}_{\rm mid}(r=\xi)=2k_{\rm mid},\quad V^{\prime\prime}_{\rm out}(r=\alpha)=2k_{\rm out},$ (11)
where $V^{\prime\prime}(r)=d^{2}V(r)/dr^{2}$. The conditions for bistability are, therefore, given by
$k_{\rm in}>0,\qquad k_{\rm mid}<0,\qquad k_{\rm out}>0.$ (12)
Continuity of $V^{\prime}(r)$ at $r=r_{\pm}$ gives explicit formulas for the matching radii:
$r_{-}$ $=\xi\sqrt{\frac{-k_{\rm mid}}{k_{\rm in}-k_{\rm mid}}},\qquad r_{+}=\sqrt{\frac{k_{\rm out}\alpha^{2}-k_{\rm mid}\xi^{2}}{\,k_{\rm out}-k_{\rm mid}\,}}.$ (13)
For the sign choice above, these roots are real and positive, satisfying $0<r_{-}<\xi<r_{+}<\alpha$ for appropriate parameter sets.
Continuity of $V(r)$ at $r_{\pm}$ fixes $C_{\rm mid}$ and $C_{\rm out}$:
$C_{\rm mid}$ $=\frac{k_{\rm in}-k_{\rm mid}}{2}r_{-}^{2}-b_{\rm mid}\ln r_{-}+C_{\rm in},$ (14)
$C_{\rm out}$ $=\frac{k_{\rm mid}-k_{\rm out}}{2}r_{+}^{2}+(b_{\rm mid}-b_{\rm out})\ln r_{+}+C_{\rm mid}.$ (15)
One may set $C_{\rm in}=0$ without loss of generality.
The potential $V(r)$ in Eq. (10) has two minima at $r=0$ and $r=\alpha$. Then, the values of $V(r)$ at the these minima are, respectively, given by $V(0)=0$, and
$V(\alpha)$ $=\frac{k_{\rm in}}{2}r_{-}^{2}+\frac{k_{\rm mid}}{2}\left(r_{+}^{2}-r_{-}^{2}\right)+\frac{k_{\rm out}}{2}\left(\alpha^{2}-r_{+}^{2}\right)+b_{\rm mid}\ln\!\left(\frac{r_{+}}{r_{-}}\right)+b_{\rm out}\ln\!\left(\frac{\alpha}{r_{+}}\right)$
$=\frac{1}{2}k_{\rm out}\alpha^{2}\ln\frac{k_{\rm out}\alpha^{2}-k_{\rm mid}\xi^{2}}{\alpha^{2}(k_{\rm out}-k_{\rm mid})}+\frac{1}{2}k_{\rm mid}\xi^{2}\ln\left(\frac{k_{\rm out}-k_{\rm mid}}{k_{\rm in}-k_{\rm mid}}\frac{-k_{\rm mid}\xi^{2}}{k_{\rm out}\alpha^{2}-k_{\rm mid}\xi^{2}}\right),$ (16)
where we have used Eqs. (9), (13), (14), and (15). We can consider the cases for (i) $V(\alpha)<0$, (ii) $V(\alpha)=0$, and (iii) $V(\alpha)>0$. If we choose $k_{\mathrm{in}}=k_{\mathrm{out}}=\xi=1$, the marginal line for $V(\alpha)=0$ is plotted as shown in Fig. 2.
Figure 2: Phase diagram of the sign of $V(\alpha)$ in the parameter space $(-k_{\mathrm{mid}},\alpha)$ for $k_{\mathrm{in}}=k_{\mathrm{out}}=\xi=1$, where the solid line represents the boundary defined by $V(\alpha)=0$.
II.3 Map to Schrödinger equation
II.3.1 Framework
We can map from Fokker-Planck equation to Schrödinger equation. This might help us to analyze the model. Let us introduce $\mathbb{P}(\bm{r},t)$ defined as
$\mathbb{P}(r,t):=e^{\beta V(r)/2}P(r,t),$ (17)
where $\mathbb{P}(\bm{r},t)$ satisfies
$\frac{\partial}{\partial t}\mathbb{P}(r,t)$ $=-\mathbb{H}\mathbb{P}(r,t),$ (18)
$\mathbb{H}$ $:=-e^{\beta V(r)/2}\mathbb{L}e^{-\beta V(r)/2}=-T\nabla^{2}+V_{S}(r).$ (19)
Recall that the Schrödinger mapping gives
$V_{S}(r):=\frac{1}{4T}\,(V^{\prime}(r))^{2}+\frac{1}{2r}\frac{d}{dr}\!\big(rV^{\prime}(r)\big).$ (20)
For the potential in Eq. (9), $V_{S}(r)$ is given by
$V_{S}(r)=\begin{cases}k_{\rm in}+\dfrac{k_{\rm in}^{2}}{4T}r^{2},&0\leq r<r_{-},\\[10.00002pt] k_{\rm mid}+\dfrac{k_{\rm mid}b_{\rm mid}}{2T}+\dfrac{k_{\rm mid}^{2}}{4T}r^{2}+\dfrac{b_{\rm mid}^{2}}{4Tr^{2}},&r_{-}\leq r<r_{+},\\[10.00002pt] k_{\rm out}+\dfrac{k_{\rm out}b_{\rm out}}{2T}+\dfrac{k_{\rm out}^{2}}{4T}r^{2}+\dfrac{b_{\rm out}^{2}}{4Tr^{2}},&r\geq r_{+}.\end{cases}$ (21)
Demanding extrema for $V_{S}(r)$ at $r=0,\xi$, and $\alpha$ fixes
$b_{\rm mid}=-k_{\rm mid}\,\xi^{2},\quad b_{\rm out}=-k_{\rm out}\,\alpha^{2}.$ (22)
Equation (21) can be rewritten as
$V_{S}(r)=k_{\mathrm{I}}+\frac{k_{\mathrm{I}}b_{\mathrm{I}}}{2T}+\frac{k_{\mathrm{I}}^{2}}{4T}r^{2}+\frac{b_{\mathrm{I}}^{2}}{4Tr^{2}},$ (23)
where the suffix I represents the region $\mathrm{I}\in\{\mathrm{in},\mathrm{mid},\mathrm{out}\}$ with $b_{\mathrm{in}}=0$. The advantage of using the map to Schrödinger equation is obvious because of $\mathbb{H}^{\dagger}=\mathbb{H}$.
We should note the following: i) Each segment of $V_{S}(r)$ is solvable quadratic plus inverse-square potential. ii) The solutions of the corresponding radial Schrödinger equation are expressible in terms of confluent hypergeometric functions. iii) The matching conditions at $r=r_{-}$ and $r=r_{+}$ enforce continuity of both $V(r)$ and $V^{\prime}(r)$, which automatically ensure the continuity of $V_{S}(r)$ and the physical wavefunctions. This construction provides a fully solvable model for a two-dimensional bistable potential with spherical symmetry.
Now, the problem is mapped onto the eigenvalue problem of the operator $\mathbb{H}$:
$\mathbb{H}\varphi_{m,n}(r)=\lambda_{m,0}\varphi_{m,n}(r),$ (24)
where we have used the relations of $\varphi_{m,n}(r)$ satisfying
$\ell_{m,n}(\bm{r})=e^{\beta V(r)/2}\varphi_{m,n}(r),\quad\mathrm{and}\quad r_{m,n}(r)=e^{-\beta V(r)/2}\varphi_{m,n}(r),$ (25)
where $\ell_{m,n}(\bm{r})$ and $r_{m,n}(\bm{r})$ are, respectively, the left and right eigenvectors, which satisfy
$\mathbb{L}r_{m,n}(\bm{r})$ $=-\lambda_{m,0}r_{m,n}(\bm{r}),$ (26)
$\mathbb{L}^{\dagger}\ell_{m,n}(\bm{r})$ $=\lambda_{m,0}\ell_{m,n}(\bm{r}),\quad\mathbb{L}^{\dagger}:=-\frac{\bm{r}}{r}V^{\prime}(r)\cdot\bm{\nabla}+T\nabla^{2}.$ (27)
Note that the two-dimensional Schrödinger equation has discrete eigenvalues characterized by a set of non-negative integers $(m,n)$, where the eigenvalue for an isotropic potential $V_{S}(r)$ should be independent of $n$, i.e., $\lambda_{m,n}=\lambda_{m,0}$ for an arbitrary non-negative integer $n$.
When we begin with an isotropic (equilibrium) condition Eq. (6), the solution of Schrödinger equation is independent of the second index $n$. Thus, the orthonormal condition the orthonormal relation
$\int d\bm{r}\ell_{m,n}(r)r_{m^{\prime},n^{\prime}}(r)=\delta_{mm^{\prime}}\delta_{nn^{\prime}}$ (28)
is converted into
$\int d\bm{r}\varphi_{m,0}(r)\varphi_{m^{\prime},0}(r)=\delta_{mm^{\prime}}.$ (29)
Then, we can use the spectrum decomposition of $P(r,t)$ as
$P(r,t)$ $=P_{\mathrm{eq}}(r,\beta)+e^{-\beta V(r)/2}\sum_{m=2}^{\infty}a_{m}\phi_{m}e^{-\tilde{\lambda}_{m}t}$
$\approx P_{\mathrm{eq}}(r,\beta)+e^{-\beta V(r)/2}\left\{a_{2}\phi_{2}(r)e^{-\tilde{\lambda}_{2}t}+a_{3}\phi_{3}(r)e^{-\tilde{\lambda}_{3}t}\right\},$ (30)
where
$a_{m}$ $=\frac{2\pi\delta_{n0}}{Z(\beta_{\mathrm{ini}})}\int_{0}^{\infty}drr\exp\left[\left(\frac{\beta}{2}-\beta_{\mathrm{ini}}\right)V(r)\right]\phi_{m}(r).$ (31)
For such an initial condition, we shall introduce
$\phi_{m}(r):=\varphi_{m-1,0}(r),\quad\tilde{\lambda}_{m}:=\lambda_{m-1,0}=\lambda_{m-1,n}.$ (32)
Note that $a_{m}$ corresponds to a one-dimensional counterpart, where $m=1$ and $m=2$ represent the zero mode and slowest relaxation mode, respectively. We also write the equation:
$F_{m}(\beta_{\mathrm{ini}}):=\frac{\partial a_{m}}{\partial\beta_{\mathrm{ini}}}=-2\pi\int_{0}^{\infty}drre^{\beta V(r)/2}\phi_{m}(r)[V(r)-\langle V\rangle_{\beta_{\mathrm{ini}}}]P_{\mathrm{ini}}(r)$ (33)
with $\langle V\rangle_{\beta_{\mathrm{ini}}}:=2\pi\int_{0}^{\infty}drrV(r)P_{\mathrm{ini}}(r)$, which is important to detect the necessary condition $F_{2}(\beta_{\mathrm{ini}}^{*})=0$ at $\beta_{\mathrm{ini}}=\beta_{\mathrm{ini}}^{*}$.
The approximate expression in Eq. (II.3.1) is useful in the long-time limit. Although the eigenvalues are degenerate for $n\neq 0$, the eigenvalues are non-degenerate for modes with $n=0$. Thus, the eigenvalues satisfy the order $\tilde{\lambda}_{1}=0<\tilde{\lambda}_{2}<\tilde{\lambda}_{3}<\tilde{\lambda}_{4}<\cdots$ with finite gaps between $\tilde{\lambda}_{n}$ and $\tilde{\lambda}_{n+1}$. Therefore, the late-stage dynamics is dominated by the slowest modes associated with $\tilde{\lambda}_{2}$ and $\tilde{\lambda}_{3}$,
Using Eq. (II.3.1), we can write that the late-stage relaxation dynamics of the expectation value of an arbitrary observable $\hat{A}$ is described by
$\langle\hat{A}\rangle(t)\approx 2\pi\int_{0}^{\infty}drrP_{\mathrm{eq}}(r,\beta)\hat{A}+2\pi\int_{0}^{\infty}drr\hat{A}\left[e^{-\beta V(r)/2}\{a_{2}\phi_{2}(r)e^{-\tilde{\lambda}_{2}t}+a_{3}\phi_{3}(r)e^{-\tilde{\lambda}_{3}t}\right],$ (34)
where $\langle\hat{A}\rangle(t):=2\pi\int_{0}^{\infty}drrP(r,t)\hat{A}$. Thus, the relaxation of $\langle\hat{A}\rangle(t)$ is dominated by the slow modes, proportional to $a_{2}$ and $a_{3}$.
Note that 1D model can be solved using a set of harmonic and anti-harmonic potentials [54]. The idea of using harmonic and anti-harmonic potentials to describe tunneling in an asymmetric double-well potential in quantum mechanics was introduced by Dekker [60] and later developed by Song [61, 62]. Our analysis shares similarities with previous papers [60, 61, 62], the target is completely different from them.
II.3.2 Condition to observe the Mpemba effect
Let us consider Eqs. (II.3.1) and (34) in the large $t$ regime, where the relaxation term proportional to $a_{2}$ is only relevant. Then, the previous papers [4, 51, 53, 54] indicated that the condition to observe the Mpemba effect is that $a_{2}$ has a maximum or a minimum against $\beta_{\mathrm{ini}}$. Indeed, this condition is necessary because the peak of $a_{2}$ at $\beta_{\mathrm{ini}}^{*}$ corresponds to the slowest initial temperature to approach the final equilibrium state after the quench. This suggests that two relaxation processes, one starting from $\beta_{\mathrm{ini}}^{*}$ and the other from $\beta_{\mathrm{ini}}<\beta_{\mathrm{ini}}^{*}$, may cause a crossing of a thermodynamic observable. However, this condition may be insufficient, as it does not guarantee that the slowest eigenmode crosses another relaxation mode proportional to $a_{3}$. To improve the defect of the previous condition, only $a_{2}$, we will propose a more direct crossing condition for a monotone measure, such as the KL divergence, to observe the Mpemba effect later.
II.3.3 Matching conditions
Since we adopt the connected potential Eq. (9), which can be converted into the effective potential Eq. (21), we should impose the connection conditions at $r=r_{-}$ and $r=r_{+}$. The connection conditions should read
$P_{\mathrm{in}}(r_{-};t)$ $=P_{\mathrm{mid}}(r_{-};t),\quad J_{r,\mathrm{in}}(r_{-};t)=J_{r,\mathrm{mid}}(r_{-};t)$ (35)
$P_{\mathrm{mid}}\left(r_{+};t\right)$ $=P_{\mathrm{out}}\left(r_{+};t\right),\quad J_{r,\mathrm{mid}}\left(r_{+};t\right)=J_{r,\mathrm{out}}\left(r_{+};t\right),$ (36)
where $P_{\mathrm{in}}(r;t)$, $P_{\mathrm{mid}}(r;t)$, and $P_{\mathrm{out}}(r;t)$ are $P(r,t)$ for $0\leq r<r_{-}$, $r_{-}\leq r<r_{+}$, and $r\geq r_{+}$, respectively. Similarly, $J_{r}(r,\theta;t):=\bm{e}_{r}\cdot\bm{J}(r,\theta;t)$ is the radial component of the current $\bm{J}$ introduced in Eq. (5), and $J_{r,\mathrm{in}}$, $J_{r,\mathrm{mid}}$, and $J_{r,\mathrm{out}}$ are $J_{r}$ for $0\leq r<r_{-}$, $r_{-}\leq r<r_{+}$, and $r\geq r_{+}$, respectively. Since $U^{\prime}(r)$ is continuous at the connection points, the continuity condition of the current can read the continuity condition of $\partial P(r,t)/\partial r$ if the radial current $J_{r}$ is regular at the connection points. Using Eqs. (5) and (II.3.1), the continuity conditions can read
$\phi_{n}^{\mathrm{in}}(r_{-})$ $=\phi_{n}^{\mathrm{mid}}(r_{-}),\quad\frac{d}{dr}\phi_{n}^{\mathrm{in}}(r)|_{r=r_{-}}=\frac{d}{dr}\phi_{n}^{\mathrm{mid}}(r)|_{r=r_{-}},$ (37)
$\phi_{n}^{\mathrm{mid}}(r_{+})$ $=\phi_{n}^{\mathrm{out}}\left(r_{+}\right),\quad\frac{d}{dr}\phi_{n}^{\mathrm{mid}}(r)|_{r=r_{+}}=\frac{d}{dr}\phi_{n}^{\mathrm{out}}(r)|_{r=r_{+}},$ (38)
where $\phi_{n}^{\mathrm{in}}(r)$, $\phi_{n}^{\mathrm{mid}}(r)$, and $\phi_{n}^{\mathrm{out}}(r)$ are $\phi_{n}(r)$ for $0\leq r<r_{-}$, $r_{-}\leq r<r_{+}$, and $r\geq r_{+}$, respectively.
III Mode analysis
III.1 Eigenvalue equation
Let us rewrite $\mathbb{H}$ introduced in Eq. (19) using a polar coordinate $r:=\sqrt{x^{2}+y^{2}}$ under circular symmetric cases as
$\mathbb{H}$ $=-\frac{T}{r}\frac{d}{dr}\left(r\frac{d}{dr}\right)+V_{S}(r).$ (39)
Thus, we can rewrite Eq. (24) as
$\phi_{n}^{\prime\prime}(r)+\frac{\phi_{n}^{\prime}(r)}{r}-\left(\beta V_{S}(r)-\beta\tilde{\lambda}_{n}\right)\phi_{n}(r)=0.$ (40)
where $\tilde{\lambda}_{n}:=\lambda_{n-1,0}$ with a positive integer $n=1,2,3,\cdots$. Let us introduce the auxiliary parameters
$\nu_{\mathrm{I}}:=\frac{|b_{\mathrm{I}}|}{2}\sqrt{\frac{\beta}{T}},\qquad\gamma_{\mathrm{I}}:=\frac{\beta|k_{\mathrm{I}}|}{2}.$ (41)
In this paper, we take the positive branch $\nu=|\nu|$ (choose the sign consistent with regularity at $r=0$) and $\gamma\geq 0$.
Introducing
$z_{\mathrm{I}}:=\gamma_{\mathrm{I}}r^{2},\qquad\Phi_{\mathrm{I},n}(z):=\phi_{\mathrm{I},n}(r)r^{-\nu_{\mathrm{I}}}e^{z_{\mathrm{I}}/2},$ (42)
where $\Phi_{\mathrm{I},n}(z_{\mathrm{I}})$ satisfies Kummer’s confluent-hypergeometric equation (see Appendix A)
$z\Phi^{\prime\prime}_{\mathrm{I},n}(z)+(1+\nu_{\mathrm{I}}-z)\Phi^{\prime}_{\mathrm{I},n}(z)-\mu_{\mathrm{I},n}\Phi_{\mathrm{I},n}(z)=0,$ (43)
with the parameter
$\mu_{\mathrm{I},n}:=\frac{1+\nu_{\mathrm{I}}}{2}+\frac{\beta\left(\lambda_{n}-k_{\mathrm{I}}-\tfrac{k_{\mathrm{I}}|b_{\mathrm{I}}|}{2T}\right)}{4\gamma_{\mathrm{I}}}.$ (44)
Two independent solutions of the confluent-hypergeometric equation are $M(\mu;1+\nu;z):={}_{1}F_{1}(\mu;1+\nu;z)$ and $U(\mu;1+\nu;z)$ defined in Eq. (79) (see Appendix A for the definitions of them and their properties). Hence, the general solution of Eq. (40) on the region can be expressed as
$\phi_{\mathrm{I},n}(r)=r^{\nu_{\mathrm{I}}}e^{-\tfrac{\gamma_{\mathrm{I}}r^{2}}{2}}\left[A_{\mathrm{I}}M\left(\mu_{\mathrm{I},n};1+\nu_{\mathrm{I}};\gamma_{\mathrm{I}}r^{2}\right)+B_{\mathrm{I}}U\left(\mu_{\mathrm{I},n};1+\nu_{\mathrm{I}};\gamma_{\mathrm{I}}r^{2}\right)\right],$ (45)
with constants $A_{\mathrm{I}}$ and $B_{\mathrm{I}}$ to be fixed by boundary/matching conditions.
III.2 Solution of Schrödinger equation
First, we consider this situation. Because a solution should not diverge at $r\to 0$ and $r\to\infty$, one gets
$\phi_{n}(r)=\begin{cases}A_{\mathrm{in}}\Phi_{\mathrm{in,M}}^{(1)}(r)&(0\leq r<r_{-})\\ A_{\mathrm{mid}}\Phi_{\mathrm{mid,M}}^{(1)}(r)+B_{\mathrm{mid}}\Phi_{\mathrm{mid,U}}^{(1)}(r)&(r_{-}\leq r<r_{+})\\ B_{\mathrm{out}}\Phi_{\mathrm{out,U}}^{(1)}(r)&(r\geq r_{+})\end{cases},$ (46)
where $\Phi_{\mathrm{I,M}}^{(1)}(r)$ and $\Phi_{\mathrm{I,U}}^{(1)}(r)$ satisfy Eq. (43), which are, respectively, given by
$\Phi_{\mathrm{I},\mathrm{M}}^{(1)}(r):=r^{\nu_{\mathrm{I}}}e^{-\frac{\gamma_{\mathrm{I}}r^{2}}{2}}M(\mu_{\mathrm{I}};1+\nu_{\mathrm{I}};\gamma_{\mathrm{I}}r^{2}),\quad\Phi_{\mathrm{I},\mathrm{U}}^{(1)}(r):=r^{\nu_{\mathrm{I}}}e^{-\frac{\gamma_{\mathrm{I}}r^{2}}{2}}U(\mu_{\mathrm{I}};1+\nu_{\mathrm{I}};\gamma_{\mathrm{I}}r^{2})$ (47)
with Eq. (79).
Using this solution, the linear matching conditions at $r=r_{-}$ and $r=r_{+}$ given by Eqs. (37) and (38) become
$\begin{pmatrix}\Phi^{(1)}_{\rm in,M}(r_{-})&-\Phi^{(1)}_{\rm mid,M}(r_{-})&-\Phi^{(1)}_{\rm mid,U}(r_{-})&0\\ \Phi^{(2)}_{\rm in,M}(r_{-})&-\Phi^{(2)}_{\rm mid,M}(r_{-})&-\Phi^{(2)}_{\rm mid,U}(r_{-})&0\\ 0&\Phi^{(1)}_{\rm mid,M}(r_{+})&\Phi^{(1)}_{\rm mid,U}(r_{+})&-\Phi^{(1)}_{\rm out,U}(r_{+})\\ 0&\Phi^{(2)}_{\rm mid,M}(r_{+})&\Phi^{(2)}_{\rm mid,U}(r_{+})&-\Phi^{(2)}_{\rm out,U}(r_{+})\end{pmatrix}\begin{pmatrix}A_{\mathrm{in}}\\ A_{\mathrm{mid}}\\ B_{\mathrm{mid}}\\ B_{\mathrm{out}}\end{pmatrix}=\begin{pmatrix}0\\ 0\\ 0\\ 0\end{pmatrix}$. (48)
where $\Phi_{\mathrm{I},\mathrm{M}}^{(2)}(r)$ and $\Phi_{\mathrm{I},\mathrm{U}}^{(2)}(r)$ are given by
$\Phi_{\mathrm{I},\mathrm{M}}^{(2)}(r):=\frac{\partial}{\partial r}\Phi_{\mathrm{I},\mathrm{M}}^{(1)}(r),\quad\Phi_{\mathrm{I},\mathrm{U}}^{(2)}(r):=\frac{\partial}{\partial r}\Phi_{\mathrm{I},\mathrm{U}}^{(1)}(r).$ (49)
Nontrivial solutions exist only if
$\det\mathcal{M}(\lambda)=0,$ (50)
where the matrix $\mathcal{M}$ is defined as
$\mathcal{M}:=\begin{pmatrix}\Phi^{(1)}_{\rm in,M}(r_{-})&-\Phi^{(1)}_{\rm mid,M}(r_{-})&-\Phi^{(1)}_{\rm mid,U}(r_{-})&0\\ \Phi^{(2)}_{\rm in,M}(r_{-})&-\Phi^{(2)}_{\rm mid,M}(r_{-})&-\Phi^{(2)}_{\rm mid,U}(r_{-})&0\\ 0&\Phi^{(1)}_{\rm mid,M}(r_{+})&\Phi^{(1)}_{\rm mid,U}(r_{+})&-\Phi^{(1)}_{\rm out,U}(r_{+})\\ 0&\Phi^{(2)}_{\rm mid,M}(r_{+})&\Phi^{(2)}_{\rm mid,U}(r_{+})&-\Phi^{(2)}_{\rm out,U}(r_{+})\end{pmatrix}$. (51)
Here, we can rewrite Eq. (50) explicitly as
$\det\mathcal{M}(\lambda)$ $=\left[\Phi^{(1)}_{\rm in,M}(r_{-})\Phi^{(2)}_{\rm mid,M}(r_{-})-\Phi^{(2)}_{\rm in,M}(r_{-})\Phi^{(1)}_{\rm mid,U}(r_{-})\right]\left[\Phi^{(1)}_{\rm mid,U}(r_{+})\Phi^{(2)}_{\rm out,U}(r_{+})-\Phi^{(2)}_{\rm mid,U}(r_{+})\Phi^{(1)}_{\rm out,U}(r_{+})\right]$
$\hskip 10.00002pt+\left[\Phi^{(1)}_{\rm in,M}(r_{-})\Phi^{(2)}_{\rm mid,U}(r_{-})-\Phi^{(2)}_{\rm in,M}(r_{-})\Phi^{(1)}_{\rm mid,U}(r_{-})\right]\left[\Phi^{(2)}_{\rm mid,M}(r_{+})\Phi^{(1)}_{\rm out,U}(r_{+})-\Phi^{(1)}_{\rm mid,M}(r_{+})\Phi^{(2)}_{\rm out,U}(r_{+})\right].$ (52)
Interestingly, the solutions of Eq. (50) are classified into two types: trivial solutions with integer eigenvalues and nontrivial solutions. As will be seen, only the nontrivial solutions are relevant.
IV Demonstration of the Mpemba effect
In this section, we examine whether the Mpemba effect can be observed. First, we examine whether $a_{2}$ has a minimum or a maximum against $\beta_{\mathrm{ini}}$, which is regarded as a necessary condition to observe the Mpemba effect. We also show the behavior of $a_{m}$ for $m\geq 3$ against $\beta_{\mathrm{ini}}$.
IV.1 (i) The case for $V(\alpha)<0$
IV.1.1 Eigenvalues and eigenfunctions
Figure 3: Plots of $V(r)$ and $V_{S}(r)$ as functions of $r$ for the set of parameters in Eq. (53).
First, we examine the case for $V(\alpha)<0$. Thus, we adopt parameters
$T=\beta=1,\ k_{\mathrm{in}}=1.0,\ k_{\mathrm{mid}}=-0.5,\ k_{\mathrm{out}}=0.8,\ \xi=1.0,\ \alpha=3.0.$ (53)
Substituting these parameters into Eqs. (13) and (22), one gets
$\ b_{\mathrm{mid}}$ $=0.5,\ b_{\mathrm{out}}=-7.2,\ r_{-}=\frac{1}{\sqrt{3}},\ r_{+}=\sqrt{\frac{7.7}{1.3}}\approx 2.4337372338,$ (54)
$\mu_{\mathrm{in},n}$ $=\tilde{\lambda}_{n},\ \mu_{\mathrm{mid},n}=\tilde{\lambda}_{n}+\frac{5}{4},\ \mu_{\mathrm{out},n}=\frac{5}{8}\tilde{\lambda}_{n}+\frac{18}{5}$ (55)
Figure 3 shows $V(r)$ and $V_{S}(r)$ for Eq. (53).
Figure 4 represents the values of $\det\mathcal{M}(\lambda)$ against $\lambda$ using Eqs. (50) and (53) for $T=1$, $k_{\mathrm{in}}=1.0$, $k_{\mathrm{mid}}=-0.5$, $k_{\mathrm{out}}=0.8$, $\xi=1.0$, and $\alpha=3.0$. It exhibits various solutions of $\det(\mathcal{M}(\lambda))=0$. Interestingly, odd crossing points with $\partial_{\lambda}\det(\mathcal{M}(\lambda))>0$ are expressed as integers, i.e., $\lambda=0,1,2,3,4,\cdots$. However, these zeros cannot be the eigenvalues of Eq. (40) because they satisfy another condition
$\mu_{\mathrm{I,n}}=-k_{\mathrm{I}}\quad(k_{\mathrm{I}}\in\mathbb{Z}_{\geq 0}),$ (56)
which leads to overcomplete conditions in Eq. (50).
Figure 4: Plot of $\det\mathcal{M}(\lambda)$ against $\lambda$ using Eq. (53) for $T=1$, $k_{\mathrm{in}}=1.0$, $k_{\mathrm{mid}}=-0.5$, $k_{\mathrm{out}}=0.8$, $\xi=1.0$, and $\alpha=3.0$.
Taking into account the normalization $\int d\bm{r}\phi_{m}(r)^{2}=1$ for arbitrary non-negative integer $m$, i.e.,
$\int_{0}^{\infty}drr\phi_{m}(r)^{2}=\frac{1}{2\pi},$ (57)
we obtain the eigenvalues $\tilde{\lambda}_{m}$ and the coefficients $(A_{\mathrm{in},m},\ A_{\mathrm{mid},m},\ B_{\mathrm{mid},m},\ B_{\mathrm{out},m})$ for $2\leq m\leq 7$ as listed in Table 1. Substituting these parameters into Eq. (46), we can determine the eigenfunction explicitly.
Table 1: Sets of $\tilde{\lambda}_{m},\ A_{\mathrm{in},m},\ A_{\mathrm{mid},m},\ B_{\mathrm{mid},m},\ B_{\mathrm{out},m}$ for $m=2,3,4,5,6,7$.
$m$ $\tilde{\lambda}_{m}$ $A_{\mathrm{in},m}$ $A_{\mathrm{mid},m}$ $B_{\mathrm{mid},m}$ $B_{\mathrm{out},m}$
22 $0.20249779968052614$ $0.24497167652225776$ $0.318031788080577$ 0.001519418698544472-0.001519418698544472 $0.012462026556778565$
33 $1.72785595359095$ $0.2543519205519021$ $0.31138740327823805$ $0.005785139374360521$ 0.01665270948531903-0.01665270948531903
44 $3.0371972138891414$ $0.319072354782721$ $0.5820667697217128$ $0.019666161001434822$ $0.008894179895148529$
55 $4.488629801287407$ $0.3102828007442467$ $0.3803646569368832$ 0.000380533758631454-0.000380533758631454 0.003309913507067175-0.003309913507067175
66 $6.0311394594$ $0.3224708012$ $0.8548841062$ 0.0003441691-0.0003441691 $0.0008296590$
77 $7.5563667569$ $0.3491231142$ $0.4119514102$ $0.0000021085$ 0.0001693156-0.0001693156
IV.1.2 $a_{m}$ for $m=2,3,4,5$
Figure 5(a) shows $\beta_{\mathrm{ini}}$ dependence of $a_{2}$ and Fig. 5(b) shows $a_{3}$, $a_{4}$, and $a_{5}$. It confirms the existence of a peak of $a_{2}$ against $\beta_{\mathrm{ini}}$, which is the necessary condition to observe the Mpemba effect. Remarkably, we expect that the overdamped Langevin system confined in a two-dimensional asymmetric potential exhibits the peak of $a_{2}$ without the introduction of a hard wall. This result contrasts with that in the one-dimensional asymmetric potential, where the Mpemba effect disappears if the system contains no hard walls [53, 54].
Figure 5: Plots of (a) $a_{2}(\beta_{\mathrm{ini}},\beta)$ and (b) $a_{m}(\beta_{\mathrm{ini}},\beta)$ ($m=3,4,5$) against the inverse temperature $\beta_{\mathrm{ini}}$ for $T=1$, $k_{\mathrm{in}}=1.0$, $k_{\mathrm{mid}}=-0.5$, $k_{\mathrm{out}}=0.8$, $\xi=1.0$, and $\alpha=3.0$.
IV.2 Higher-order contributions in the spectrum decomposition
So far, we have assumed that the time evolution of the distribution function can be described by the lowest few eigenmodes in Eq. (II.3.1). This is true if the Mpemba effect is dominated by the late-stage dynamics, but the initial relaxation might be important.
To examine the assumption that the initial relaxation is irrelevant, we examine
$P(r,t)$ $=P_{\mathrm{eq}}(r,\beta)+e^{-\beta V(r)/2}\sum_{m=2}^{N_{\mathrm{tr}}}a_{m}\phi_{m}e^{-\tilde{\lambda}_{m}t},$ (58)
where $N_{\mathrm{tr}}$ is the truncation of the spectrum decomposition. The initial distribution $P_{\mathrm{ini}}$ can be approximated as
$P(r,0)$ $=P_{\mathrm{eq}}(r,\beta)+e^{-\beta V(r)/2}\sum_{m=2}^{N_{\mathrm{tr}}}a_{m}\phi_{m}.$ (59)
Figure 6: (a) Plots of $P_{\mathrm{ini}}(r)$ at $\beta_{\mathrm{ini}}=0.82$, given by Eqs. (6) and  (59) for $N_{\mathrm{tr}}=3,5,7$. (b) Time evolution of the difference of the distribution $P(r,t)-P_{\mathrm{eq}}(r,\beta)$ for $N_{\mathrm{tr}}=7$, where we adopt Eq. (53) for the set of parameters.
Figure 6(a) shows the initial distribution $P_{\mathrm{ini}}$, q. (59), for $\beta_{\mathrm{ini}}=0.82$ with various $N_{\mathrm{tr}}(=3,5,7)$. Although the approximation is not accurate for $N_{\mathrm{tr}}=3$, it becomes sufficiently accurate when $N_{\mathrm{tr}}=7$. Figure 6(b) plots the time evolution of the distribution function for $N_{\mathrm{tr}}=7$. For clarity, we plot the difference from the final state $P_{\mathrm{eq}}(r,\beta)$. Since the smallest eigenvalue is $\tilde{\lambda}_{2}=0.20249779968052614$, the distribution function is found to converge to $P_{\mathrm{eq}}(r,\beta)$ for $t\gg 1/\tilde{\lambda}_{2}\approx 5$. Furthermore, as shown in Appendix C, the Mpemba effect is still observed in terms of the difference in the KL divergence even when higher-order modes are included. Therefore, Eq. (II.3.1) provides a valid description of the dynamics.
IV.3 Condition of the crossing of the KL divergence
Although we have confirmed the existence of a peak of $a_{2}$ against $\beta_{\mathrm{ini}}$, this may not be sufficient to observe the Mpemba effect. To observe the Mpemba effect, we need a crossing condition for the observable. In this subsection, we clarify the sufficient condition to observe the Mpemba effect.
Among many possible observables, we choose the KL divergence defined as
$D^{\mathrm{KL}}(P(\beta,\beta_{\mathrm{ini}};t)||P_{\mathrm{eq}}(\beta))$ $:=\int d\bm{r}P(\bm{r},t)\ln\frac{P(\bm{r},t)}{P_{\mathrm{eq}}(r,\beta)}=2\pi\sum_{n=2}^{\infty}\frac{(-1)^{n}}{n(n-1)}\int_{0}^{\infty}drr\frac{(\delta P)^{n}}{P_{\mathrm{eq}}^{n-1}},$ (60)
which is a typical monotone measure, where $\delta P:=P(\bm{r},t)-P_{\mathrm{eq}}(\bm{r})$. If we are interested in the late stage of relaxation dynamics or the deviation from the initial state and the final state is small, $D^{\mathrm{KL}}(P(t)||P_{\mathrm{eq}})$ can be approximated as
$D^{\mathrm{KL}}(P(\beta,\beta_{\mathrm{ini}};t)||P_{\mathrm{eq}}(\beta))$ $=2\pi\int_{0}^{\infty}drr\frac{(\delta P)^{2}}{2P_{\mathrm{eq}}}+O(\delta P^{3})\approx\frac{1}{2}\sum_{n\geq 2}C_{n}\left(a_{n}\right)^{2}e^{-2\tilde{\lambda}_{n}t},$ (61)
where we have introduced
$C_{n}:=\|\phi_{n}\|^{2}_{P_{\rm eq}^{-1}}>0$ (62)
with $\|\phi_{n}\|^{2}_{P_{\rm eq}^{-1}}:=2\pi Z(\beta)\int_{0}^{\infty}drr\phi_{n}(r)^{2}$. Here, $C_{n}$ should be independent of $n$ because each mode $\phi_{n}(r)$ satisfies the normalization. Thus, we denote $C_{n}$ as $C^{*}$.
If we are interested in the late-stage dynamics, the modes associated with the slowest two modes $\tilde{\lambda}_{2}$ and $\tilde{\lambda}_{3}$ are dominant because $e^{-\tilde{\lambda}_{m}t}\ll 1$ for $m\geq 4$ is negligible. Therefore, we adopt the two-mode approximation as
$D^{\mathrm{KL}}(P(t)||P_{\mathrm{eq}})\approx\frac{C^{*}}{2}(a_{2}^{2}e^{-2\tilde{\lambda}_{2}t}+a_{3}^{2}e^{-2\tilde{\lambda}_{3}t}).$ (63)
Thus, the condition of a crossing of the KL divergence, characterized by Eq. (63), is expressed as
$\mathcal{F}(\beta_{\mathrm{ini}}^{\sharp},\beta_{\mathrm{ini}}^{*}):=\frac{\Delta_{3}}{\Delta_{2}}<-1,$ (64)
where
$\Delta_{n}:=\left(a_{n}^{(\mathrm{A})}\right)^{2}-\left(a_{n}^{(\mathrm{B})}\right)^{2}.$ (65)
for two different initial conditions A and B. See Appendix B for the derivation of Eq. (64). This quadratic form is valid only when $P(r,t)$ is sufficiently close to equilibrium; in particular, it does not reproduce the exact value at $t=0$ when the initial condition is another equilibrium distribution. Therefore, the condition Eq. (64) should be regarded as a necessary (but not always sufficient) condition for crossing within the validity of the truncated expansion.
The peak of $a_{2}$ against $\beta_{\mathrm{ini}}$ is located around $\beta_{\mathrm{ini}}^{*}\approx 0.862$. Let us choose another initial condition:
$\beta_{\mathrm{ini}}^{\sharp}=0.500.$ (66)
In this case, we evaluate
$a_{2}(\beta_{\mathrm{ini}}^{\sharp})$ $\approx 0.0692734,\ a_{2}(\beta_{\mathrm{ini}}^{*})\approx 0.07230438,$
$a_{3}(\beta_{\mathrm{ini}}^{\sharp})$ $\approx-0.07144855,\ a_{3}(\beta_{\mathrm{ini}}^{*})\approx-0.07098009.$ (67)
Then, we obtain
$\Delta_{2}$ $:=a_{2}(\beta_{\mathrm{ini}}^{\sharp})^{2}-a_{2}(\beta_{\mathrm{ini}}^{\mathrm{*}})^{2}\approx-0.00424624<0,$ (68)
$\Delta_{3}$ $:=a_{3}(\beta_{\mathrm{ini}}^{\sharp})^{2}-a_{3}(\beta_{\mathrm{ini}}^{\mathrm{*}})^{2}\approx 0.00130880>0.$ (69)
Figure 7: The demonstration of the Mpemba effect through the time evolution of the difference of the KL divergence $\Delta D:=D^{\mathrm{KL}}(P(0.70,1;t)||P_{\mathrm{eq}}(0.70))-D^{\mathrm{KL}}(P(0.82,1;t)||P_{\mathrm{eq}}(0.82))$ using Eq. (63) (solid line) and Eq. (93) (dotted line), where we use Eq. (53) for the set of parameters. We also plot the dashed line, which includes the higher-order contributions as in Eq. (58) with $N_{\mathrm{tr}}=7$.
We plot the time evolution of the KL divergence $\Delta D(\beta^{\sharp},\beta^{*}):=D^{\mathrm{KL}}(P(\beta^{\sharp},\beta;t)||P_{\mathrm{eq}}(\beta^{\sharp}))-D^{\mathrm{KL}}(P(\beta^{*},\beta;t)||P_{\mathrm{eq}}(\beta))$ during a cooling process starting from $\beta_{\mathrm{ini}}^{\sharp}=0.70$ and $\beta_{\mathrm{ini}}^{*}=0.82$ to $\beta=1$ in Fig. 7 (see the solid line). It is apparent to observe the crossing around $t\approx 1$, which exhibits the Mpemba effect. We now compare the KL divergence, including higher-order terms, by using the expansion of $P(r,t)$ truncated at $N_{\mathrm{tr}}=7$ as the dashed line. The dotted line represents the semi-analytic result of Eq. (93) (see Appendix C). From these results, the two-mode approximation works well, although the crossing time depends on $N_{\mathrm{tr}}$. Also, the semi-analytic result Eq. (93) is close to the numerical result with $N_{\mathrm{tr}}=7$ around the crossing time. Therefore, we conclude that (i) the crossing condition Eq. (64) can be used for the detection of the crossing of the KL divergence, and (ii) the crossing time can be evaluated by the semi-analytic result Eq. (93).
We draw the phase diagram to clarify the region in which the Mpemba effect (shaded region) can be observed in Fig. 8 for the set of parameters in Eq. (53), when we fix one initial temperature as $\beta_{\mathrm{ini}}=\beta_{\mathrm{ini}}^{*}$. Remarkably, the Mpemba effect cannot be observed if the two initial temperatures are too close, and it occurs for
$0.66919732\lesssim\beta_{\mathrm{ini}}^{\sharp}\lesssim 0.71013378.$ (70)
Figure 8: Plot of $\mathcal{F}(\beta_{\mathrm{ini}}^{\sharp},\beta_{\mathrm{ini}}^{*}):=\Delta_{2}/\Delta_{3}$ against the initial temperature $\beta_{\mathrm{ini}}^{\sharp}$, where one initial temperature is fixed at $\beta^{*}_{\mathrm{ini}}=0.82$, and use Eq. (53) for the set of parameters. The shaded region indicates where the Mpemba effect can be observed as in Eq. (64).
IV.4 Case (ii) $V(\alpha)=0$
Figure 9: Plots of (a) $a_{2}(\beta_{\mathrm{ini}},\beta)$ and (b) $a_{m}(\beta_{\mathrm{ini}},\beta)$ ($m=3,4,5$) against the inverse temperature $\beta_{\mathrm{ini}}$ for $T=1$, $k_{\mathrm{in}}=-k_{\mathrm{mid}}=k_{\mathrm{out}}=1$, $\xi=1.0$, and $\alpha=1.84485714$.
Figure 10: Plot of $\mathcal{F}(\beta_{\mathrm{ini}}^{\sharp},\beta_{\mathrm{ini}}^{*}):=\Delta_{2}/\Delta_{3}$ against the initial temperature $\beta_{\mathrm{ini}}^{\sharp}$, where one initial temperature is fixed at $\beta^{*}_{\mathrm{ini}}=1.15$. The parameter set given in Eq. (71) is used.
Next, let us consider case (ii), i.e., $V(\alpha)=0$. In this case, the well depth at $r=0$ is equal to that at $r=\alpha$. Fixing the parameters
$T=1,\ k_{\mathrm{in}}=-k_{\mathrm{mid}}=k_{\mathrm{out}}=1,\ \xi=1.0,\ \alpha=1.84485714,$ (71)
we plot Fig. 9 as the $\beta_{\mathrm{ini}}$ dependence of $a_{m}$ ($m=2,3,4,5$). In contrast to case (i), a peak of $a_{2}$ exists at $\beta_{\mathrm{ini}}=\beta_{\mathrm{ini}}^{*}\approx 1.15>1$. Therefore, if the Mpemba-like effect exists, it should be the inverse Mpemba effect in the heating process.
Interestingly, despite the existence of a peak in $a_{2}$, we do not observe the Mpemba effect for $V(\alpha)=0$. Indeed, the condition Eq. (64) is violated in this case, as shown in Fig. 10, in which $\mathcal{F}(\beta^{\sharp},\beta^{*})>-1$.
Now, let us examine whether the inverse Mpemba effect can be observed for $V(\alpha)=0$ by controlling the curvature of the potential. Figure 11 shows the value of $\mathcal{F}(\beta_{\mathrm{ini}}^{\sharp},\beta_{\mathrm{ini}}^{*}=1.15)$ against $\beta_{\mathrm{ini}}^{\sharp}$. Throughout the entire range of $\beta_{\mathrm{ini}}^{\sharp}$, $\mathcal{F}$ remains larger than 1-1, indicating that the Mpemba effect cannot exist. This is an interesting example, in which the inverse Mpemba effect indeed cannot be observed even if $a_{2}$ has a peak.
Figure 11: Phase diagram of the Mpemba effect when we change $k_{\mathrm{mid}}$ and $k_{\mathrm{out}}$ by fixing $k_{\mathrm{in}}=\xi=1$ under the condition $V(\alpha)=0$. Here, the crosses indicate that the Mpemba effect is not observed.
IV.5 Case (iii) $V(\alpha)>0$
Next, let us consider case (iii) $V(\alpha)>0$. Here, we adopt parameters
$T=1,\ k_{\mathrm{in}}=k_{\mathrm{out}}=1.5,\ \xi=1.0,\ -k_{\mathrm{mid}}=\alpha=1.5.$ (72)
In this case, $a_{2}(\beta_{\mathrm{ini}},\beta)$ decreases with $\beta_{\mathrm{ini}}$ monotonically as shown in Fig. 12(a) while the behaviors of $a_{m}(\beta_{\mathrm{ini}},\beta)$ for $m=2,3,4$ are the similar to the previous cases.
Figure 12: Plots of (a) $a_{2}(\beta_{\mathrm{ini}},\beta)$ and (b) $a_{m}(\beta_{\mathrm{ini}},\beta)$ ($m=3,4,5$) against the inverse temperature $\beta_{\mathrm{ini}}$ when we choose the set of parameters Eq. (72).
IV.6 Phase diagram of the Mpemba effect
Here, we draw the phase diagram to distinguish the region of the occurrence of the Mpemba effect from the non-Mpemba region under the control of a set of parameters. Figure 13 exhibits a phase diagram on the plane of $(k_{\mathrm{mid}},\alpha)$ by fixing $k_{\mathrm{in}}=k_{\mathrm{out}}=\xi=1$, where the circles and crosses correspond to the occurrence of the Mpemba effect and absence of the Mpemba effect, respectively. We also plot open triangles where $a_{2}$ has a peak, while Eq. (64) is not satisfied. Interestingly, the open triangles are located not only in the vicinity of $V(\alpha)=0$ but also in the region with large $\alpha$. We need to clarify the mechanism for the appearance of such regions.
Figure 13: Phase diagram of the Mpemba effect when we change $k_{\mathrm{mid}}$ and $\alpha$ by fixing $k_{\mathrm{in}}=k_{\mathrm{out}}=\xi=1$. The open circles and crosses indicate regions where the Mpemba effect can be observed and cannot be observed, respectively. The open triangles represent that $a_{2}$ exhibits a peak while Eq. (64) is not satisfied. The crosses show that $a_{2}$ has no peak. The solid line represents $V(\alpha)=0$ as shown in Fig. 2.
V Discussion
We now discuss our results, focusing on the difference between 1D and 2D, and thermomajorization.
V.1 Difference between 1D and 2D
Remarkably, the Mpemba effect can be observed even in an unbounded system in a two-dimensional situation, while we need to introduce a wall to confine the particle in a one-dimensional asymmetric potential [53, 54]. The difference may be from the forbidden region for $r<0$ in 2D cases, i.e., there is an effective wall at $r=0$ in the 2D case. More explicitly, the equilibrium condition for 2D system is expressed as
$P_{\mathrm{eq}}(r)\propto re^{-\beta V(r)}.$ (73)
Thus, the effective potential mapping onto 1D systems can be written as
$V_{\mathrm{eff}}(r)=V(r)-T\ln r.$ (74)
This effective potential prohibits the location of a test particle for $r<0$, which can be regarded as an effective wall at $r=0$.
The necessary condition to observe the Mpemba effect is whether
$F_{2}(\beta_{\mathrm{ini}}):=\frac{\partial a_{2}}{\partial\beta_{\mathrm{ini}}}=-2\pi\int_{0}^{\infty}drre^{\beta V(r)/2}\phi_{2}(r)[V(r)-\langle V\rangle_{\beta_{\mathrm{ini}}}]P_{\mathrm{ini}}(r)$ (75)
has, at least, one zero for $\beta_{\mathrm{ini}}=\beta_{\mathrm{ini}}^{*}$. We demonstrate that $F_{2}(\beta_{\mathrm{ini}})$ has a different sign in the low-temperature and high-temperature limits for the case (i) $V(\alpha)<V(0)=0$, as shown in Appendix D.1. Thus, we conclude the existence $F_{2}(\beta_{\mathrm{ini}}^{*})=0$ at a $\beta_{\mathrm{ini}}=\beta_{\mathrm{ini}}^{*}$. Mathematically, the difference in the Jacobian between 2D and 1D leads to different conclusions, while $r=0$ can be regarded as an effective potential in 1D problem. Interestingly, we can prove the absence of positive $\beta_{\mathrm{ini}}^{*}$ to satisfy $F_{2}(\beta_{\mathrm{ini}}^{*})$ for the case (iii) $V(\alpha)>0$ (see Appendix D.2). When we regard $r=0$ as an effective potential in 1D system, this result is analogous to that in 1D case with a wall on one side [53, 54].
V.2 Thermomajorization
Vu and Hayakawa [56] proposed the thermomajorization for a general condition to observe the Mpemba effect in arbitrary monotone measures. This concept is useful as a general criterion to observe the Mpemba effect, but we may not need the thermomajorization once we know the expansion coefficients $a_{2}$, $a_{3}$, and the others explicitly. Indeed, the argument in Sec. IV.3 can be generalized to an arbitrary observable $\hat{A}$ by using Eq. (34), where the existence of a maximum or a minimum in $a_{2}$ is still the necessary condition to observe the Mpemba effect for an arbitrary observable. If we know $a_{2}$ and $a_{3}$ explicitly, one can obtain the crossing condition corresponding to Eq. (64) for an arbitrary observable $\hat{A}$. Thus, we do not have to worry about the possibility that some monotone measure may not cross during the relaxation, while the KL divergence exhibits the crossing.
VI Conclusion
By introducing an exactly solvable model, we have obtained exact eigenmodes of the Fokker-Planck equation in a bistable potential. From spectral analysis, we identify the necessary and sufficient conditions for observing the Mpemba effect. Using the slowest and the second slowest eigenmodes of the particle’s motion in a two-dimensional potential, starting from an equilibrium initial condition after a quench into another equilibrium state, we demonstrate the occurrence of a crossing of the KL-divergence, if $V(\alpha)<V(r_{\mathrm{min}})=0$ where $r_{\mathrm{min}}$ is the location of the minimum of $V(r)$ near $r=0$. In contrast to the 1D case, we can observe the Mpemba effect for a particle constrained in a 2D double potential without the introduction of any confined walls because $r=0$ can be regarded as an effective wall. Interestingly, $a_{2}$ has a peak, but neither the Mpemba effect nor the inverse Mpemba effect cannot be observed for $V(\alpha)=0$. For $V(\alpha)>0$, there is no peak for $a_{2}$ and the Mpemba effect cannot be observed.
Acknowledgment
We thank Y. Liu for stimulating discussion. HH also thanks N. Ohga, F. van Wijland, R. Chetrite, and T. Van Vu for fruitful discussion. This work is partially supported by JSPS Kakenhi (Grant Nos. JP24K06974, JP24K07193, and JP25K01063).
Appendix A Properties of Tricomi’s $U$-function
The Tricomi (or Tricomi–Confluent) function $U(a,b,z)$ is the second, linearly independent solution of Kummer’s (confluent-hypergeometric) differential equation
$zy^{\prime\prime}(z)+(b-z)y^{\prime}(z)-ay(z)=0.$ (76)
The other standard solution is Kummer’s $M$-function (also denoted $_{1}F_{1}$[63]:
$M(a,b,z):={}_{1}F_{1}(a;b;z)=\sum_{k=0}^{\infty}\frac{(a)_{k}}{(b)_{k}}\frac{z^{k}}{k!},$ (77)
where $(q)_{k}:=\Gamma(q+k)/\Gamma(q)$ with the Gamma function $\Gamma(x):=\int_{0}^{\infty}dtt^{x-1}e^{-t}$ is the Pochhammer symbol.
A convenient and frequently used representation of $U(a,b,z)$ is the integral formula (valid for $\Re a>0,\ \Re z>0$):
$U(a,b,z):=\frac{1}{\Gamma(a)}\int_{0}^{\infty}e^{-zt}t^{a-1}(1+t)^{b-a-1}dt.$ (78)
By analytic continuation, this defines $U(a,b,z)$ for general complex parameters and $z$ (with the usual branch cuts).
The connection formula expressing $U$ in terms of $M$ is
$U(a,b,z)=\frac{\Gamma(1-b)}{\Gamma(a-b+1)}M(a,b,z)+\frac{\Gamma(b-1)}{\Gamma(a)}z^{1-b}M(a-b+1,2-b,z),$ (79)
which is valid by analytic continuation for non-integer $b$ and extends to integer $b$ by limit. Useful identities and asymptotics are presented as
$\frac{d}{dz}M(a,b,z)=\frac{a}{b}M(a+1,b+1,z),$ (80)
$\frac{d}{dz}U(a,b,z)=-a\,U(a+1,b+1,z),$ (81)
$U(a,b,z)\sim z^{-a}\Big(1+O(1/z)\Big)\qquad(\lvert z\rvert\to\infty,\ \Re z>0).$ (82)
In our application (with the notation of the main text), one sets
$a=\mu,\qquad b=1+\nu,\qquad z=\gamma r^{2},$ (83)
so that $U(\mu;1+\nu;\gamma r^{2})$ denotes Tricomi’s function with those parameters. The identity (81) is particularly useful when computing derivatives of the radial solution
$\phi(r)=r^{\nu}e^{-\tfrac{\gamma r^{2}}{2}}U(\mu,1+\nu,\gamma r^{2})$ (84)
appearing in the matching conditions.
Appendix B Quantitative Criterion for Crossing of the KL Divergence
In this appendix, we derive a simple quantitative criterion for the crossing of the KL divergence between two different initial temperatures.
Consider two initial states AA and $B$. Introducing
$\Delta D(t):=D_{A}(t)-D_{B}(t),$ (85)
with the aid of Eq. (61), we write
$\Delta D(t)=\frac{C^{*}}{2}\sum_{n\geq 1}\Delta_{n}e^{-2\lambda_{n}t}.$ (86)
A crossing occurs if there exists a time $t^{\ast}>0$ such that
$\Delta D(t^{\ast})=0.$ (87)
In the minimal two-mode approximation ($n=2,3$), this gives
$\Delta_{2}e^{-2\tilde{\lambda}_{2}t^{\ast}}+\Delta_{3}e^{-2\tilde{\lambda}_{3}t^{\ast}}=0.$ (88)
Solving for $t^{\ast}$,
$t^{\ast}=\frac{1}{2(\tilde{\lambda}_{3}-\tilde{\lambda}_{2})}\ln\left(\frac{-\Delta_{3}}{\Delta_{2}}\right).$ (89)
A real positive solution exists if and only if Eq. (64) is satisfied.
Equation (64) shows that a non-monotonic dependence of $a_{2}$ on the initial temperature is necessary but not sufficient for the Mpemba effect defined via KL crossing. The amplitude contrast must be strong enough to reverse the ordering relative to faster modes.
Appendix C Contributions of higher-order terms
In this appendix, we examine how the results change if we include higher-order terms for $\Delta D(\beta^{\sharp},\beta^{*}):=D^{\mathrm{KL}}(P(\beta^{\sharp},\beta;t)||P_{\mathrm{eq}}(\beta^{\sharp}))-D^{\mathrm{KL}}(P(\beta^{*},\beta;t)||P_{\mathrm{eq}}(\beta))$. Let us consider higher-order terms in the KL divergence as
$D^{\mathrm{KL}}(P(\beta,\beta_{\mathrm{ini}};t)||P_{\mathrm{eq}}(\beta))$ $=\frac{1}{2}\sum_{n\geq 2}C^{*}(a_{n})^{2}e^{-2\tilde{\lambda}_{n}t}-\frac{1}{6}\sum_{\ell\geq 2}\sum_{m\geq 2}\sum_{n\geq 2}D_{\ell mn}a_{\ell}a_{m}a_{n}e^{-(\tilde{\lambda}_{\ell}+\tilde{\lambda}_{m}+\tilde{\lambda}_{n})t}$
$\hskip 10.00002pt+\frac{1}{12}\sum_{k\geq 2}\sum_{\ell\geq 2}\sum_{m\geq 2}\sum_{n\geq 2}E_{k\ell mn}a_{k}a_{\ell}a_{m}a_{n}e^{-(\tilde{\lambda}_{k}+\tilde{\lambda}_{\ell}+\tilde{\lambda}_{m}+\tilde{\lambda}_{n})t}+O(\delta P^{5}),$ (90)
with $D_{\ell mn}:=2\pi Z(\beta)^{2}\int_{0}^{\infty}drre^{\beta V(r)/2}\phi_{\ell}(r)\phi_{m}(r)\phi_{n}(r)$ and $E_{k\ell mn}:=2\pi Z(\beta)^{3}\int_{0}^{\infty}drre^{\beta V(r)}\phi_{k}(r)\phi_{\ell}(r)\phi_{m}(r)\phi_{n}(r)$. The first few terms of $D_{\ell mn}$ and $E_{k\ell mn}$ are given by
$D_{222}$ $\approx 7.34748352,\ D_{223}\approx 2.52984461,\ D_{233}\approx 3.31794833,\ D_{333}\approx-0.78230974,$ (91)
$E_{2222}$ $\approx 10.71271826,\ E_{2223}\approx 5.09076249,\ E_{2233}\approx 4.20693782,\ E_{2333}\approx 2.40306438,\ E_{3333}\approx 3.84968712.$ (92)
Since $\tilde{\lambda}_{n}(\geq 0)$ increases with increasing $n$, we here employ the same approximation as used in Eq. (63), keeping only $\tilde{\lambda}_{2}$ and $\tilde{\lambda}_{3}$. Therefore, Eq. (90) becomes approximately
$D^{\mathrm{KL}}(P(\beta,\beta_{\mathrm{ini}};t)||P_{\mathrm{eq}}(\beta))=D^{(2)}(t)+D^{(3)}(t)+D^{(4)}(t)+\cdots$ (93)
where
$D^{(2)}(t):$ $=\frac{1}{2}C^{*}\left(a_{2}^{2}e^{-2\tilde{\lambda}_{2}t}+a_{3}^{2}e^{-2\tilde{\lambda}_{3}t}\right),$ (94)
$D^{(3)}(t):$ $=-\frac{1}{6}\left(D_{222}a_{2}^{3}e^{-3\tilde{\lambda}_{2}t}+3D_{223}a_{2}^{2}a_{3}e^{-(2\tilde{\lambda}_{2}+\tilde{\lambda}_{3})t}+3D_{233}a_{2}a_{3}^{2}e^{-(\tilde{\lambda}_{2}+2\tilde{\lambda}_{3})t}+D_{333}a_{3}^{3}e^{-3\tilde{\lambda}_{3}t}\right),$ (95)
$D^{(4)}(t):$ $=\frac{1}{12}\left(E_{2222}a_{2}^{4}e^{-4\tilde{\lambda}_{2}t}+4E_{2223}a_{2}^{3}a_{3}e^{-(3\tilde{\lambda}_{2}+\tilde{\lambda}_{3})t}+6E_{2233}a_{2}^{2}a_{3}^{2}e^{-2(\tilde{\lambda}_{2}+\tilde{\lambda}_{3})t}\right$. (96)
$\left.\qquad+4E_{2333}a_{2}a_{3}^{3}e^{-(\tilde{\lambda}_{2}+3\tilde{\lambda}_{3})t}+E_{3333}a_{3}^{4}e^{-4\tilde{\lambda}_{3}t}\right).$ (97)
Figure 7 contains the dotted line as the results of Eq. (93) and the dashed line as the numerical result with $N_{\mathrm{tr}}=7$, where we adopt Eq. (63) for the set of parameters. Even if we include the higher-order contributions, the prediction of the Mpemba effect by using Eq. (64) is not destroyed. Therefore, the Mpemba effect obtained in the low-order approximation persists even when higher-order contributions are taken into account.
The expansion of the KL divergence in terms of eigenmodes is useful to describe the relaxation behavior for $t>0$. However, for the present problem, the initial condition is an equilibrium distribution at $\beta_{\mathrm{ini}}$ [see Eq. (6)]. Therefore, the KL divergence at $t=0$ can be evaluated exactly from its definition Eq. (60), without relying on any mode expansion.
Equation  (60) at $t=0$ can read
$D^{\mathrm{KL}}(0)=\int drP_{\mathrm{eq}}(r,\beta_{\mathrm{ini}})\ln\frac{P_{\mathrm{eq}}(r,\beta_{\mathrm{ini}})}{P_{\mathrm{eq}}(r,\beta)}=(\beta-\beta_{\mathrm{ini}})\langle V\rangle_{\beta_{\mathrm{ini}}}+\ln\frac{Z(\beta)}{Z(\beta_{\mathrm{ini}})},$ (98)
where we have used Eq. (6). This is an exact expression for $\Delta D(0)$.
Thus, the difference of the KL divergences for two initial temperatures $A$ and $B$ at $t=0$ is given by
$\Delta D(0)=D^{\mathrm{KL}}_{A}(0)-D^{\mathrm{KL}}_{B}(0),$ (99)
which can be evaluated exactly from equilibrium thermodynamics.
On the other hand, for $t>0$, the deviation from equilibrium can be expressed using the spectral decomposition Eq. (II.3.1), which leads to
$D^{\mathrm{KL}}(t)=\frac{1}{2}\sum_{n\geq 2}C_{n}(\hat{a}_{n})^{2}e^{-2\tilde{\lambda}_{n}t}+O(e^{-3\tilde{\lambda}_{2}t}).$ (100)
Therefore, the time-dependent difference becomes
$\Delta D(t)=\Delta D(0)+\frac{1}{2}\sum_{n\geq 2}C_{n}\Delta_{n}\left(e^{-2\tilde{\lambda}_{n}t}-1\right)+O(e^{-3\tilde{\lambda}_{2}t}),$ (101)
where $\Delta_{n}:=(\hat{a}_{n}^{(A)})^{2}-(\hat{a}_{n}^{(B)})^{2}$.
A crossing requires that $\Delta D(t)$ changes sign for some $t>0$, i.e.,
$\Delta D(0)\Delta D(t\to\infty)<0.$ (102)
Since $e^{-2\tilde{\lambda}_{n}t}-1<0$, the relaxation contributions tend to reduce the initial difference. Therefore, a necessary condition for crossing is that the modal contributions overcome the initial gap:
$\sum_{n\geq 2}C_{n}\Delta_{n}<0\quad(\text{if }\Delta D(0)>0),$ (103)
and similarly for the opposite case.
In the late-stage regime, keeping only the two slowest modes, we obtain
$\Delta D(t)\simeq\Delta D(0)+\frac{C^{*}}{2}\left[\Delta_{2}(e^{-2\tilde{\lambda}_{2}t}-1)+\Delta_{3}(e^{-2\tilde{\lambda}_{3}t}-1)\right].$ (104)
This expression provides a crossing condition when we consider many eigenmodes.
Thus, the role of higher-order modes is not to determine the initial value, but to modify how the system departs from the exact initial KL divergence. Their contribution can shift the crossing time quantitatively, but the qualitative mechanism of the crossing remains controlled by the competition between $\Delta D(0)$ and the slowest relaxation modes.
Appendix D Sign change of $F_{2}(\beta_{\rm ini})$ for the 2D double-well potential
In this appendix, we prove that for the two-dimensional double-well potential defined in Eq. (9), $F_{2}(\beta_{\mathrm{ini}})$ introduced Eq. (75) changes sign between the limits $\beta_{\rm ini}\to 0$ and $\beta_{\rm ini}\to\infty$. Let us rewrite $F_{2}(\beta_{\mathrm{ini}})$ as
$F_{2}(\beta_{\rm ini})=-2\pi\int_{0}^{\infty}drr\phi_{2}(r)\left[V(r)-\langle V\rangle_{\beta_{\rm ini}}\right]\frac{e^{-(\beta_{\rm ini}-\beta/2)V(r)}}{Z(\beta_{\rm ini})}.$ (105)
D.1 (i) The case for $V(\alpha)<V(0)$
Hete, we assume $V(\alpha)<V(0)$ as the case (i).
D.1.1 High-temperature limit $\beta_{\rm ini}\to 0$
In the high-temperature limit, the initial distribution becomes broad and the measure is dominated by the large-volume region. Because of the Jacobian factor $r$, the dominant contribution comes from the vicinity of the potential minima at $r=\alpha$ rather than from the barrier region near $r=0$.
For the first radial excited mode,
$\phi_{2}(r)\begin{cases}<0,&r\approx 0,\\ >0,&r\approx\alpha.\end{cases}$ (106)
In the wells,
$V(\alpha)<\langle V\rangle_{\beta_{\rm ini}\to 0}.$ (107)
Hence,
$\phi_{2}(r)\,\big(V(r)-\langle V\rangle_{\beta_{\mathrm{ini}}}\big)<0\quad\text{in the dominant region}.$ (108)
Because of the overall minus sign in Eq. (105), we conclude
$F_{2}(0)>0.$ (109)
Low-temperature limit $\beta_{\rm ini}\to\infty$
As $\beta_{\rm ini}\to\infty$, the measure concentrates near $r=\alpha$. We therefore introduce the local coordinate
$x:=r-\alpha.$ (110)
The potential near the global minimum is expressed as
$V(r)=V(\alpha)+\frac{k}{2}x^{2},$ (111)
where $k=a_{\mathrm{out}}$.
Expand the eigenfunction:
$\phi_{2}(r)=c_{0}+c_{1}x+\frac{a_{2}}{2}x^{2}+O(x^{3}),$ (112)
where
$c_{0}:=\phi_{2}(\alpha).$ (113)
Introduce the scaled variable
$y:=\sqrt{\beta_{\rm ini}k}x,$ (114)
we can write
$e^{-\beta_{\rm ini}V(r)}=e^{-\beta_{\rm ini}V(\alpha)}\exp\!\left(-\frac{y^{2}}{2}\right)\left[1+O(\beta_{\rm ini}^{-1/2})\right].$ (115)
Using Eq. (111), we immediately obtain
$V(r)-\langle V\rangle_{\beta_{\rm ini}}=\frac{k}{2}x^{2}+O(x^{3}).$ (116)
Next, we use the expansion
$r\phi_{2}(r)$ $=(\alpha+x)\left(c_{0}+c_{1}x+\frac{c_{2}}{2}x^{2}\right)+O(x^{3})$ (117)
$=\alpha c_{0}+(\alpha c_{1}+c_{0})x+\left(\frac{\alpha c_{2}}{2}+c_{1}\right)x^{2}+O(x^{3}).$ (118)
Multiplying Eq. (116) with Eq. (117), we obtain
$r\phi_{2}(r)\bigl(V(r)-\langle V\rangle_{\beta_{\mathrm{ini}}}\bigr)$ $=\frac{k}{2}\left[\alpha c_{0}x^{2}+(\alpha c_{1}+c_{0})x^{3}+O(x^{4})\right].$ (119)
Under the Gaussian limit, we write
$\langle x^{2}\rangle_{\beta_{\mathrm{ini}}}=\frac{1}{\beta_{\rm ini}k},\qquad\langle x^{3}\rangle_{\beta_{\mathrm{ini}}}=0,\qquad\langle x^{4}\rangle_{\beta_{\mathrm{ini}}}=\frac{3}{(\beta_{\rm ini}k)^{2}}.$ (120)
Therefore, we obtain
$\int_{0}^{\infty}drr\phi_{2}(r)\left(V(r)-\langle V\rangle_{\beta_{\mathrm{ini}}}\right)=\frac{\alpha c_{0}}{2\beta_{\rm ini}}+O(\beta_{\rm ini}^{-3/2}).$ (121)
We, thus, conclude the relation
$F_{2}(\beta_{\rm ini})=-\frac{\alpha\,\phi_{2}(\alpha)}{2\beta_{\rm ini}}+O(\beta_{\rm ini}^{-3/2})<0,\qquad\beta_{\rm ini}\to\infty.$ (122)
D.1.2 Conclusion
From the above discussion, we have established
$F_{2}(0)>0,\qquad\lim_{\beta_{\mathrm{ini}}\to\infty}F_{2}(\beta_{\mathrm{ini}})<0.$ (123)
By continuity, there exists at least one $\beta_{\rm ini}^{*}$ such that
$F_{2}(\beta_{\rm ini}^{*})=0.$ (124)
Hence, the Mpemba effect necessarily occurs for the two-dimensional double-well potential.
D.2 (iii) The case $V(0)<V(\alpha)$
In this subsection, we prove the absence of the Mpemba effect for (iii) the case
$V(\alpha)>V(0)=0.$ (125)
In this regime, the global minimum of the potential is located at $r=0$, while $r=\alpha$ corresponds to a metastable minimum. Our strategy is to show $F_{2}(\beta_{\mathrm{ini}})$ introduced in Eq. (75) has a definite sign for all $\beta_{\rm ini}>0$. This excludes any extremum of $\hat{a}_{2}(\beta_{\rm ini})$, which is a necessary condition for the Mpemba effect. The argument, here, can be regarded as a customized version of the original idea in Ref. [55], which develops a microscopic theory of the Mpemba effect.
From Eq. (75) we have
$F_{2}(\beta_{\rm ini})=-\frac{2\pi}{Z(\beta_{\rm ini})}\int_{0}^{\infty}drr\phi_{2}(r)e^{-\beta_{\rm ini}V(r)}\left[V(r)-\langle V\rangle_{\beta_{\rm ini}}\right],$ (126)
where
$\langle V\rangle_{\beta_{\rm ini}}=\frac{2\pi}{Z(\beta_{\rm ini})}\int_{0}^{\infty}drrV(r)e^{-\beta_{\rm ini}V(r)}.$ (127)
Introducing the probability measure
$d\mu_{\beta_{\rm ini}}(r)=\frac{2\pi}{Z(\beta_{\rm ini})}re^{-\beta_{\rm ini}V(r)}dr,$ (128)
Eq. (126) can be rewritten as
$F_{2}(\beta_{\rm ini})=-\mathrm{Cov}_{\beta_{\rm ini}}\left(\phi_{2}(r),V(r)\right),$ (129)
where the covariance $\mathrm{Cov}_{\beta_{\rm ini}}\left(\phi_{2}(r),V(r)\right):=\phi_{2}(r)\left[V(r)-\langle V\rangle_{\beta_{\rm ini}}\right]$ is taken with respect to $d\mu_{\beta_{\rm ini}}$.
Therefore, the sign of $F_{2}$ is opposite to the covariance between $\phi_{2}(r)$ and $V(r)$ under the Gibbs measure.
For $V(\alpha)>0$, the energetic ordering of the two wells is
$V(0)<V(\alpha).$ (130)
The first nontrivial radial eigenfunction $\phi_{2}(r)$ has a single node and opposite signs in the two wells. Up to an overall normalization, we choose
[…]
Hence, the sign of $\phi_{2}$ increases when moving from the lower-energy well to the higher-energy well. In particular, for any $r_{1},r_{2}$ belonging respectively to the inner and outer wells,
$V(r_{1})<V(r_{2})\quad\Longrightarrow\quad\phi_{2}(r_{1})<\phi_{2}(r_{2}).$ (132)
Thus $\phi_{2}(r)$ is an increasing function of $V(r)$ in the sense of well-to-well ordering.
Let us decompose the integration domain into the inner-well region $\Omega_{0}$ and outer-well region $\Omega_{\alpha}$. Because of Eq. (132), $\phi_{2}(r)$ and $V(r)$ are positively ordered: larger values of $V$ are associated with larger values of $\phi_{2}$.
By Chebyshev’s integral inequality for similarly ordered functions, for any probability measure (see Appendix D)
$\mathrm{Cov}_{\beta_{\rm ini}}\left(\phi_{2}(r),V(r)\right)\geq 0.$ (133)
The inequality is strict because $\phi_{2}$ is not a constant function and the two wells have different energies. Therefore,
$\mathrm{Cov}_{\beta_{\rm ini}}\left(\phi_{2}(r),V(r)\right)>0\quad\text{for all }\beta_{\rm ini}>0.$ (134)
Importantly, the Gibbs weight $e^{-\beta_{\rm ini}V(r)}$ changes continuously with $\beta_{\rm ini}$, but does not alter the energetic ordering between the two wells. Hence, the positive ordering between $\phi_{2}$ and $V$ is preserved for all temperatures.
Combining Eqs. (129) and (134), we obtain
$F_{2}(\beta_{\rm ini})<0\quad\text{for all }\beta_{\rm ini}>0.$ (135)
Therefore $\hat{a}_{2}(\beta_{\rm ini})$ is strictly decreasing and has no extremum.
Since an extremum of $\hat{a}_{2}$ is a necessary condition for the crossing of the KL divergence, the Mpemba effect cannot occur in the regime $V(\alpha)>0$.
Chebyshev’s integral inequality for similarly ordered functions
In this subsection, we state and prove Chebyshev’s integral inequality in a form appropriate for probability measures. The proof is elementary and relies on a symmetric representation of the covariance.
Setting.
Let $(X,\mathcal{F},\mu)$ be a probability space, i.e. $\mu(X)=1$. Let $f,g:X\to\mathbb{R}$ be square-integrable functions.
We say that $f$ and $g$ are similarly ordered if for all $x,y\in X$,
$(f(x)-f(y))(g(x)-g(y))\geq 0.$ (136)
In other words, whenever $f$ increases, $g$ also increases, and whenever $f$ decreases, $g$ also decreases.
Theorem (Chebyshev’s integral inequality).
If $f$ and $g$ are similarly ordered, then
$\int_{X}f(x)g(x)d\mu(x)\geq\left(\int_{X}f(x)d\mu(x)\right)\left(\int_{X}g(x)d\mu(x)\right).$ (137)
Equivalently,
$\mathrm{Cov}_{\mu}(f,g)\geq 0.$ (138)
If the inequality in (136) is strict on a set of positive $\mu\times\mu$ measure, then the covariance is strictly positive.
Proof.
Recall that the covariance can be written symmetrically as
$\mathrm{Cov}_{\mu}(f,g)=\frac{1}{2}\int_{X}\int_{X}(f(x)-f(y))(g(x)-g(y))d\mu(x)d\mu(y).$ (139)
To verify (139), expand the integrand:
$(f(x)-f(y))(g(x)-g(y))$
$=f(x)g(x)+f(y)g(y)-f(x)g(y)-f(y)g(x).$ (140)
Integrating over $X\times X$ and using $\mu(X)=1$, we obtain
$\int_{X}\int_{X}f(x)g(x)d\mu(x)d\mu(y)$ $=\int_{X}f(x)g(x)d\mu(x),$ (141)
$\int_{X}\int_{X}f(x)g(y)d\mu(x)d\mu(y)$ $=\left(\int_{X}fd\mu\right)\left(\int_{X}gd\mu\right),$ (142)
and similarly for the remaining terms. Collecting all contributions yields
$\int_{X}\int_{X}(f(x)-f(y))(g(x)-g(y))d\mu(x)d\mu(y)=2\mathrm{Cov}_{\mu}(f,g),$ (143)
which proves (139).
Now assume that $f$ and $g$ are similarly ordered, i.e. (136) holds for all $x,y$. Then the integrand in (139) is nonnegative everywhere. Therefore,
$\mathrm{Cov}_{\mu}(f,g)\geq 0.$ (144)
If the inequality in (136) is strict on a set of positive $\mu\times\mu$ measure, then the double integral is strictly positive, and hence
$\mathrm{Cov}_{\mu}(f,g)>0.$ (145)
This completes the proof. $\square$